OpenStudy (thisisallido):

Help! I'll fan and medal! Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Two divided by x squared. + 3

OpenStudy (hyperlunatic):

what school do you go to?

OpenStudy (thisisallido):

Homeschool

OpenStudy (hyperlunatic):

do you use K12 virtual acadamey materials?

OpenStudy (hyperlunatic):

Ok are there answer choices, i want to make sure this answer is in range...

OpenStudy (thisisallido):

??? And no it's an essay question

OpenStudy (hyperlunatic):

OH its an essay question?

OpenStudy (thisisallido):

Yes unfortunately

OpenStudy (hyperlunatic):

\[y= \frac{ 2 }{ x ^{2} } + 3\]

OpenStudy (hyperlunatic):

That is the answer, can you give any instructions for the essay or do you even have them?

OpenStudy (thisisallido):

That's all it gave

OpenStudy (hyperlunatic):

ok... I really don't know abou the essay part, I can't help you there sorry. :(

OpenStudy (phi):

f( g(x) ) means replace any "x" in your equation with g(x) for example *IF* you had at the rule: \[ f(x) = \frac{2}{x} +3 \] then f( g(x) ) would be: \[ f( g(x))= \frac{2}{g(x)} +3 \] what should you make g(x) be ? i.e. what "rule" so that if you replace g(x) with its "rule" you get \[ f(g(x)) = \frac{2}{x^2} +3 \]

OpenStudy (thisisallido):

Is that what the answer is? @phi

OpenStudy (thisisallido):

@phi

OpenStudy (thisisallido):

@calculusxy

OpenStudy (phi):

it's almost the answer

OpenStudy (thisisallido):

I'm confused at what to do, that didn't make sense to me @phi

OpenStudy (phi):

The first thing is to get this idea: "x" is a "placeholder", and it is something that shows where you "put something" Think of the equation (for example) f(x)= x+3 as a "pattern" or string of letters and numbers. But everywhere there is the symbol "x", you can replace it with a "new" letter or number or string. normally, we use numbers, and we use the syntax f(3) to mean: replace x with 3 in f(x)=x+3 we would get: f(3) = 3+3

OpenStudy (phi):

another example g(x)= x+x+x what is g(1) ? we start with g(x)= x+x+x we replace x with 1: g(1) = 1+1+1

OpenStudy (phi):

once that idea makes sense, and you can do problems like f(x)= x^2 +2x+1 f(2) = 2^2 +2*2+1 (which we can then simplify using "order of operations" and arithmetic) we can tackle your problem f( g(x) )

OpenStudy (phi):

Let's say f(x) = x+1 what is f( g(x) ) ? we do the same thing as before: we replace the symbol x with g(x) : f(x) = x+1 becomes f( g(x) ) = g(x) + 1 in other words, everywhere you see "x" you erased it, and put in "g(x)" the process is simple. Of course, we don't get a number: g(x)+1 looks a bit strange, but that is ok.

OpenStudy (phi):

but what if we had the equation g(x) = x^2 + 2x + 1 ?? g(x) is the short-name for the expression x^2 + 2x + 1 if we were told f(x) = g(x) we would know f(x) = x^2 + 2x + 1 or if we knew f(x) = g(x) +1 then we know f(x) = x^2 + 2x + 1 + 1

OpenStudy (phi):

now your problem: Find f(x) and g(x) so the function can be expressed as y = f(g(x)). \[ y= \frac{2}{x^2} +3 \] we have to find f(x) and g(x) so that f( g(x) ) is that expression We can do lots of things (there are lots of ways to choose f(x)) but let's do this one: \[ f(x) = \frac{2}{x} + 3 \] now what happens if we replace x with x^2 ? we would get \[ y= \frac{2}{x^2} +3 \] which is what we want. we know (I hope) that if we use that f(x), then f(g(x)) is \[ f(g(x)) = \frac{2}{g(x)} + 3 \] and if it turned out g(x)= x^2, we would have our answer. ok, then let g(x)= x^2 we have f(x) and g(x) so that f(g(x)) is what we want